OpenStudy (anonymous):

partial decomp of (-x^2-7x+27)/X(X^2+9)

OpenStudy (abb0t):

\[\frac{A}{x}+\frac{Bx+C}{x^2+9}=\frac{-x^2-7+27}{x(x^2+9)}\]

OpenStudy (alekos):

this looks nasty!

OpenStudy (alekos):

could it also be A/x + B/(x+3) + C/(x-3) ?

OpenStudy (abb0t):

No. You don't have difference of squares in the denominator.

OpenStudy (alekos):

ahh yes youre right

OpenStudy (anonymous):

I got what he first got but now im stuck

OpenStudy (alekos):

multiply left and right sides by x(x^2+9)

OpenStudy (abb0t):

A(x\(^2\)+9)+B(x)=-x\(^2\)-7x+27 \(\Rightarrow\) use the distributive property to distribute and combine like terms

OpenStudy (anonymous):

a good start would be to let \(x=0\) and get \(9A=27\) so at least you know \(A\) right away

OpenStudy (anonymous):

so x is 3

OpenStudy (abb0t):

Yes sir.

OpenStudy (anonymous):

how do I get c

OpenStudy (anonymous):

no it is not \(x=3\) but rather \(A=3\)

OpenStudy (anonymous):

i think there maybe a mistake above

OpenStudy (anonymous):

\[\frac{A}{x}+\frac{Bx+C}{x^2+9}=\frac{-x^2-7+27}{x(x^2+9)}\] \[x(Bx+C)+A(x^2+9)=-x^2-7x+27\]

OpenStudy (anonymous):

put \(x=0\) get \[9A=27\] and so \(A=3\) giving \[x(Bx+C)+3x^2+27=-x^2-7x+27\]

OpenStudy (anonymous):

expand on the left and get \[Bx^2+Cx+3x^2+27=-x^2-7x+27\] or \[(B+3)x^2+Cx+27=-x^2-7x+27\] which gives you \(C=-7\) right away

OpenStudy (anonymous):

also tells you \((B+3)x^2=-x^2\) and so \(B+3=-1\) making \(B=-4\) lets check it

OpenStudy (anonymous):

\[\frac{A}{x}+\frac{Bx+C}{x^2+9}=\frac{-x^2-7+27}{x(x^2+9)}\] \[\frac{3}{x}+\frac{-4x-7}{x^2+9}=\frac{-x^2-7+27}{x(x^2+9)}\]